# Learning Visual Conditioning Tokens to Correct Domain Shift for Fully Test-time Adaptation

### v1

@@ II Related Work and Major Contributions
Test-time adaptation (TTA) aims to adapt a pre-trained source model to unlabeled data with domain shift during testing. It has two major approaches: test-time training [16, 25, 26] and fully test-time adaptation [17, 18, 27]. [16] proposes the first test-time training (TTT) method which updates the feature extractor network parameters according to a self-supervised loss on a proxy learning task. The TTT++ method [25] introduces a feature alignment strategy based on online moment matching. The TTT-MAE method [26] is a recent extension of TTT that utilizes the transformer backbone and replaces the self-supervision with masked autoencoders [44]. Note that all of these TTT methods need special training in the source domain.
On the other hand, fully test-time adaptation methods fine-tune the pre-trained model during inference without any access to the source data. The TENT method [17] first proposes the fully test-time adaptation by fine-tuning the BN layers. The NHL method [21] learns effective early layer representations in an unsupervised manner based on neurobiology-inspired Hebbian learning. Instead of updating the parameters of the network model, some approaches update the model inputs. DDA [45] projects the input target data onto the source domain based on a diffusion model during testing. The method proposed by [37] updates the target inputs by learning image pixel-level visual prompts for target domains with the source model parameters being frozen during testing. It has been noted that existing online model updating methods often suffer from performance degradation from sample imbalance and small batch size. EATA [19] achieves considerable performance by pre-collecting a set of in-distribution target domain samples to compute the Fisher importance for regularization. It may be infeasible in fully test-time adaptation setting. To address this issue, SAR [27] proposes to eliminate partially noisy samples with high gradients and perform flattening of model weights towards a minimum, thereby enhancing the robustness of the model. DELTA [46] enhances the estimated normalization statistics using moving averaged statistics to correct the normalized features. In this paper, we follow the same problem setting to perform online fully test-time adaptation.
@@
Recently, prompt learning has been successfully applied to vision-language models [48, 49, 50, 51, 52, 53, 54, 55]. CoOp [48] tunes the prompt of the text encoder in CLIP [56]. CoCoOp [49] learns the prompt conditioned on the input data of the model to address the out-of-distribution issue. TPT [51] optimizes the prompt of the text-encoder of CLIP during the test time by minimizing the entropy with confidence selection to improve the generalization performance. DAPL [52] learns a prompt consisting of domain-agnostic context, domain-specific context, and class label for the text-encoder with a contrastive objective to disentangle the semantic and domain representation. While these prompt methods are text-encoder-based in Vision-Language Models, it’s essential to highlight that our approach focuses solely on the vision transformer (ViT) encoder.
Besides text prompt learning, methods have been developed to learn visual prompts for computer vision tasks. The method proposed by [37] learns domain-specific prompts and domain-agnostic prompts which are attached to the target input images per pixel in continuous TTA tasks. BlackVIP [57] learns the corresponding prompt for each individual image with the same shape as the input image through a neural network without knowledge about the pre-trained model architectures and parameters. Visual Prompt Tuning (VPT) [36] prepended task-specific learnable parameters into the input sequence of each ViT encoder layer while freezing the entire pre-trained transformer encoder backbone during downstream training. This method has also been applied in the transfer learning of image synthesis [58]. The DePT method [59] expanded the VPT to test-time adaptation with memory bank-based online pseudo-labeling. However, the VPT method introduces hundreds of new tokens. It significantly increases the computational complexity of the self-attention mechanism as the computational complexity of self-attention is quadratic to the number of input tokens. All of these prompt tuning methods introduced additional learnable parameters. Different from VPT with numerous extra tokens for prompt learning, we choose to adapt the existing one class token in the transformer encoder pre-trained on the source for test-time adaptation.
@@ IV Experiments
Implementation Details
Following the official implementations of SAR111SAR: https://github.com/mr-eggplant/SAR, we use the ViT-B/16 backbone for all experiments unless explicitly stated otherwise. We use the pre-trained model weights from the timm repository [72]. For Office-Home, we fine-tuning the ViT-B/16 by replacing the original classifier head as the source model. For fair performance comparisons, all methods in each experimental setting share the same architecture and the same pre-trained model parameters. We use the SGD optimizer with the Sharpness Aware Minimization [73]. The batch size is set to 64 for all experiments except the setting Batch size = 1. The learning rates of the long-term domain-specific token and the short-term instance-specific token are set to 0.005 and 0.01, respectively. All of our experimental results are the mean values obtained by 3 times running with different random seeds, and the standard deviation is also reported on the last row of tables. All models are trained and tested on a single NVIDIA RTX3090 GPU.

### v2

@@ II Related Work and Major Contributions
Test-time adaptation aims to online adapt the trained model while testing the input samples in the target domain. [44] proposed Test-time Training (TTT) by optimizing a self-supervised loss through a proxy task on the source before adapting to the target domain. TTT++ [25] encouraged the test feature distribution to be close to the training one by matching the moments estimated online. It should be noted that this method requires specific training on the source data, which is not available in fully test-time adaptation scenarios [17]. TENT [17] fine-tuned the scale and bias parameters of batch normalization layers using an entropy minimization loss during the inference process. DUA [18] adapted the statistics of batch normalization layer only on a tiny fraction of test data and augmented a small batch of target data to adapt the model. [45] proposed a shift-agnostic weight regularization and an auxiliary task for the alignment between the source and target features. Note that this method requires the source data for computing the source prototypes. The continual test-time adaptation methods [20, 46] consider online TTA where target data is continually changing during inference. Instead of using parameters of the pre-trained model, [47] only adapted the model’s output by optimizing an objective function based on Laplacian adjusted maximum-likelihood estimation.
@@
Recently, in the realm of Vision-Language Models, several approaches have emerged that leverage prompt tuning to enhance performance [49, 50, 51, 52, 53, 54, 55, 56]. CoOp [49] introduced a method that tunes the prompt for the text encoder in CLIP [57], setting a precedent for prompt-based optimization. Extending this concept, CoCoOp [50] further conditioned the prompt learning on the input data, aiming to mitigate the out-of-distribution challenge. While these innovative methods primarily target the text encoder in Vision-Language Models, our approach stands apart by solely focusing on enhancing the vision transformer (ViT) encoder, exploring a distinct path for prompt-based learning in the visual domain..
Beyond text prompt learning, recent advancements have explored visual prompts for enhancing computer vision tasks. For instance, the method by [37] learns both domain-specific and domain-agnostic prompts that are pixel-wise attached to input images for continuous Test-Time Adaptation (TTA) tasks. BlackVIP [58] takes a different approach, learning individualized prompts for each image that share the same shape as the input, without requiring knowledge of the pre-trained model’s architecture or parameters. Visual Prompt Tuning (VPT) [36] introduced a technique that prepends task-specific learnable parameters to the input sequence of each Vision Transformer (ViT) encoder layer, while keeping the pre-trained encoder backbone frozen. This approach has been successfully applied in image synthesis tasks [59] and extended to test-time adaptation with DePT [60], leveraging a memory bank for online pseudo-labeling. However, one notable limitation of VPT is the introduction of hundreds of new tokens, which significantly escalates the computational complexity of the self-attention mechanism due to its quadratic dependency on the number of input tokens. All of these prompt tuning methods introduced additional learnable parameters. Different from VPT with numerous extra tokens for prompt learning, we choose to adapt the existing one class token in the transformer encoder pre-trained on the source for test-time adaptation.